OpenStudy (anonymous):

If p=3r is equal to 150 percent of 6r, what is the value of p/r?

OpenStudy (anonymous):

What is the question write it clearly... p/r = 3

OpenStudy (anonymous):

Since p = 3r divide both the sides by r.. you will get 3..

OpenStudy (anonymous):

oh im sorry!

OpenStudy (anonymous):

If p+3r is equal to 150 percent of 6r, what is the value of p/r?

OpenStudy (anonymous):

@waterineyes this is the question

OpenStudy (anonymous):

p + r = 150/100 of 6r \[p+r = \frac{3}{2} \times 6r \implies p+r = 9r\] \[\implies p = 8r \implies \frac{p}{r} = 8\]

OpenStudy (anonymous):

thank you again!

OpenStudy (anonymous):

Welcome dear again..